The bottom of a ladder must be placed 3feet from a wall the ladder is 12feet long how far above the ground does the ladder touch the wall
Found 2 solutions by zoomkaboom4, zurro123:
Answer by zoomkaboom4(62)

a^2+b^2= c^2 is the Pythagorean Theorem Formula
3^2+b^2= 12^2
9+b^2= 144
144-9= 135
b^2= 135
And then after you get this^ you have to square root 135, Try it!
Any more questions? Message me:)

Answer by zurro123(1)